Deployment team

→ Europe/Zurich
EVO - GridPP Deployment team meeting

EVO - GridPP Deployment team meeting

Jeremy Coles
Description
- This is the weekly DTEAM meeting - The intention is to run the meeting in EVO: http://evo.caltech.edu/evoGate/. Join the meeting in the GridPP Community area - The phone bridge number is +41 22 76 71400,. Th phone bridge ID is 124300 with code: 4880.
Minutes
    • 11:00 → 11:20
      Experiment problems/issues 20m
      Review of weekly issues by experiment/VO - LHCb - CMS - ATLAS - Other - Site problems -- APEL issues at ScotGrid-Edinburgh; IC-HEP; QMUL; RHUL; LIV-HEP; Durham (Dec) and OX-HEP -- SAM issues at IC-HEP; UCL and Manchester -- ATLAS tests at UCL; IC-HEP and LIV-HEP
    • 11:20 → 11:35
      ROC update 15m
      ops meeting update ********************* - A certificate expiry impacted APEL and some sites have reported no publishing since that time. Please could sites check if they need to republish and that they are now correctly publishing - A gLite 3.1 release update is expected this week (it is through pre-deployment and is only waiting on documentation changes) - Has anyone used the VO configurator tool? There is now a push to make it more functional but the developers need to be assured that people will use or are using it! Do we have any feedback or volunteers to look closer? There is a proposal to develop SAM tests related to the configurator (VO ID card versions used). Any comments to feed back? - There is now a regional dashboard available on the CIC portal: https://cic.gridops.org/index.php?section=roc&page=dashboard&subpage=default. Comments? ROC manager update ************************* - Next meeting is today: http://indico.cern.ch/conferenceDisplay.py?confId=23752 - CA update procedure - Getting SA1 SLDs signed - Dealing with inappropriate jobs (and reminding VOs/users of obligations) Ticket status *************** - See list attached
      more information
    • 11:35 → 11:45
      Site performance over vacation period / Q4 2007 10m
      - http://hepwww.ph.qmul.ac.uk/~lloyd/gridpp/samplots.html
    • 11:45 → 11:55
      DTEAM improvements in 2008 10m
      - Suggestions for areas of improvement - Areas team are unhappy (actions/issues/blogs/wikis/email...) - Meeting frequency and cross T2 support
    • 11:55 → 12:05
      Actions review 10m
    • 12:05 → 12:10
      AOB 5m
      - Expt. storage requirements for CCRC'08 https://twiki.cern.ch/twiki/bin/view/LCG/GSSD - There is a need to upgrade sites to latest SRM versions (only Lancaster at latest dCache 1.8.0 release and Edinburgh at latest DPM).